OpenStudy (anonymous):

Express the given product as a sum or difference containing only sines or cosines. sin(8x)cos(6x)

OpenStudy (anonymous):

Do you know the Product to Sum formula?

OpenStudy (anonymous):

Learning it right now

OpenStudy (anonymous):

ah it's sin u cos v

OpenStudy (anonymous):

Sorry this is the first problem like this I've seen

OpenStudy (anonymous):

hm maybe you didn't sub in the u and v but this would look the same. u=8x v= 6x \[\sin (u)\cos(v) = \frac{ 1 }{ 2 } (\sin(u+v) + \sin(u-v))\] After plugging in u=8x v= 6x \[\sin (8x)\cos(6x) = \frac{ 1 }{ 2 } (\sin(8x+6x) + \sin(8x-6x))\] does this look familiar?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\ = \frac{ 1 }{ 2 } (\sin(14x) + \sin(2x))\] and that's how you do it. The left side is just the original equation that we transfer from product to sum.

OpenStudy (anonymous):

Got it! thank you for explaining!

OpenStudy (anonymous):

\(\color{salmon}{\large\heartsuit}\large\text{You're Most Welcome! }\color{salmon}\heartsuit\) \(~~~~~~~~~~~~~~~\color{green}{\large\ddot\smile}\color{blue}{\large\ddot\smile}\color{salmon}{\large\ddot\smile}\color{blue}{\large\ddot\smile}\color{salmon}{\large\ddot\smile}\) \(~~~~~~~~~~~~~~~~~~~~\color{salmon}{\large\bigstar}\color{salmon}{\large\bigstar}\color{salmon}{\large\bigstar} \)

OpenStudy (anonymous):

sin(9x)cos(5x)=1/2sin(14x)+sin(4x) is that correct?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

woohoo! haha